OpenStudy (anonymous):

cos^2(pi/9) - sin^2(pi/9) Use identities to simplify. (Half angle and double angle)

OpenStudy (itsonlycdeee):

Do you know your double angle formulas?

OpenStudy (anonymous):

yes but idk how yo use it with this problem

OpenStudy (itsonlycdeee):

Oh okay. Well cos(2A) = cos^2A - sin^2A or 1 - 2sin^2 A or 2cos^2A - 1, sin (2A) = 2sinAcosA and tan (2A) = 2tanA/1-tan^2A which formula do you think matches?

OpenStudy (anonymous):

cos(2A) = cos^2A - sin^2A

OpenStudy (itsonlycdeee):

Good and your A in this case is?

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

2 times pi/9 is pi/6

OpenStudy (anonymous):

sorry got that wrong

OpenStudy (itsonlycdeee):

Actually \[\frac{ 2\pi }{ 9 }\] can't be reduced. Since it was given to you in cos^2A -sin^2A . it becomes cos (2A) , which is just cos (2*pi/9).

OpenStudy (anonymous):

ok now do u plug in>

OpenStudy (itsonlycdeee):

Well is your question looking for an answer or to just leave it as \[\cos \frac{ 2\pi }{ 9 }\] ?

OpenStudy (anonymous):

yeah it is asking for simplification. And it looks like it has been simplified. Thank you :)

OpenStudy (itsonlycdeee):

No problem! (: